34.y=x+2,
3y-2x=4
22. 4x-y=9
x-3y=16

Answer by jim_thompson5910(35256)

# 34

3y-2x=4

3(x+2)-2x=4 ... Plug in y = x+2

3x+6-2x=4

x+6=4

x=4-6

x=-2

y = x+2

y = -2+2

y = 0

So x = -2 and y = 0 giving us the ordered pair (-2, 0)
=========================================================
# 22

I'll let you tackle this one. Solve for y in the first equation. Then follow the same basic steps shown above.